# Heights and periods of algebraic cycles in families

### v1

@@ B.2. Mumford–Tate domains
The unipotent radical $V$ of $\mathbf{G}=\mathrm{MT}({\mathscr{D}})$ equals $M_{-1}\cap\mathbf{G}$ by reason of weight. Let $\mathbf{G}_{0}:=\mathbf{G}/V$ be the reductive part. Set ${\mathscr{D}}_{0}:=p({\mathscr{D}})\subseteq{\mathscr{M}}_{0}$ for the map $p$ defined in (B.6). Then ${\mathscr{D}}_{0}$ is a $\mathbf{G}_{0}({\mathbb{R}})^{+}$-orbit and is in fact a (pure) Mumford–Tate domain in the classifying space ${\mathscr{M}}_{0}$, and $\mathrm{MT}_{x_{0}}<\mathbf{G}_{0}$ for all $x_{0}\in{\mathscr{D}}_{0}$.
@@
(B.7) $p\colon\mathbf{G}\rightarrow\mathbf{G}_{0}=\mathbf{G}/V\quad\text{ and }\quad p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}.$
Fix a Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$. Identify $\mathbf{G}_{0}$ with $\{0\}\times\mathbf{G}_{0}$.
Recall that each $x_{0}\in{\mathscr{M}}_{0}$ endows $M_{-1}$ with a Hodge structure of weight $-1$. For $x_{0}\in{\mathscr{D}}_{0}$, $V$ is a sub-Hodge structure because $V$ is a $\mathbf{G}_{0}$-submodule of $M_{-1}$ and that $\mathrm{MT}_{x_{0}}<\mathbf{G}_{0}$.
Finally, consider the constant bundle $V({\mathbb{C}})\times{\mathscr{D}}_{0}\rightarrow{\mathscr{D}}_{0}$. Define the holomorphic subbundle $F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})$ to be such that the fiber over each $x_{0}\in{\mathscr{D}}_{0}$ is $F^{0}_{x_{0}}V_{{\mathbb{C}}}$.
@@
    $V({\mathbb{R}})\times{\mathscr{D}}_{0}\xrightarrow{\sim}{\mathscr{D}},\quad(v,x_{0})\mapsto\mathrm{Int}(v)\circ h_{x_{0}}.$
@@
    The complex structure on ${\mathscr{D}}$ is given by ${\mathscr{D}}=(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})$.
@@
    (B.8) $V({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq V({\mathbb{C}})\times{\mathscr{D}}_{0}\longrightarrow(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0}).$
@@
The Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ induces a bijection $\mathbf{G}({\mathbb{R}})^{+}x\cong V({\mathbb{R}})\times{\mathscr{D}}_{0}$, which is semi-algebraic. This establishes (i).
Set ${\mathscr{X}}:=\mathbf{G}({\mathbb{R}})^{+}V({\mathbb{C}})\cdot h_{x}\subseteq{\mathrm{Hom}}({\mathbb{S}}_{{\mathbb{C}}},\mathbf{G}_{{\mathbb{C}}})$, where the action is given by conjugation. The quotient $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}$ induces a natural surjective map ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and by [Pin89, 1.8(a)] each fiber of this map is a $V({\mathbb{C}})$-torsor.
The Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ induces a global section of ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and hence an isomorphism ${\mathscr{X}}\cong V({\mathbb{C}})\times{\mathscr{D}}_{0}$ over ${\mathscr{D}}_{0}$.
@@
By [Pin89, 1.8(b)], for each $x\in{\mathscr{D}}$, the fiber $\varphi^{-1}(x)$ is a principle homogeneous space under $F^{0}_{x_{0}}V_{{\mathbb{C}}}$ with $x_{0}=p(x)$. Hence $\varphi$ gives a bijection $(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})\rightarrow{\mathscr{D}}$ over ${\mathscr{D}}_{0}$. Moreover, the complex structure of ${\mathscr{D}}$ is precisely given by this bijection; see [GK24, Proof of Prop. 2.6 in Appendix A] which is a consequence of [Pea00, Thm. 3.13]. This establishes (ii).
@@
Let $\tilde{Z}_{0}\subseteq{\mathscr{D}}_{0}$ be an irreducible algebraic subset. Then for any $a\in V({\mathbb{R}})$, the subset $\{a\}\times\tilde{Z}_{0}\subseteq V({\mathbb{R}})\times{\mathscr{D}}_{0}\cong{\mathscr{D}}$ is irreducible algebraic.
@@
$\{a\}\times\tilde{Z}_{0}$ is clearly semi-algebraic in ${\mathscr{D}}$. In view of (B.8), $\{a\}\times\tilde{Z}_{0}$ is also complex analytic in ${\mathscr{D}}$. Hence, we are done. ∎
@@
Let ${\mathscr{D}}^{\prime}$ be a sub-Mumford–Tate domain of ${\mathscr{D}}$. Then under the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), we have ${\mathscr{D}}^{\prime}=(V^{\prime}({\mathbb{R}})+v_{0})\times p({\mathscr{D}}^{\prime})$ for some $v_{0}\in V({\mathbb{Q}})$.
@@
Denote by $\mathbf{G}^{\prime}=\mathrm{MT}({\mathscr{D}}^{\prime})$, $V^{\prime}:=V\cap\mathbf{G}^{\prime}$, and $\mathbf{G}^{\prime}_{0}:=\mathbf{G}^{\prime}/V^{\prime}$. Then, since ${\mathscr{D}}^{\prime}$ is a sub-Mumford–Tate domain of ${\mathscr{D}}$ and by reason of weight, $V^{\prime}$ is the unipotent radical of $\mathbf{G}^{\prime}$.
Set $\mathbf{G}^{\prime\prime}:=V\mathbf{G}^{\prime}$. Then $\mathbf{G}^{\prime\prime}$ is a subgroup of $\mathbf{G}$, with unipotent radical $V$ and reductive part $\mathbf{G}^{\prime}_{0}$. We can construct two Levi decompositions of $\mathbf{G}^{\prime\prime}$ as follows.
First, $\mathbf{G}^{\prime}_{0}=\mathbf{G}^{\prime}/V^{\prime}=\mathbf{G}^{\prime}/(V\cap\mathbf{G}^{\prime})<\mathbf{G}/V=\mathbf{G}_{0}$. Under the Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ fixed above Proposition B.3, we have $\{0\}\times\mathbf{G}^{\prime}_{0}<\mathbf{G}^{\prime\prime}$. The composite $\{0\}\times\mathbf{G}^{\prime}_{0}<\mathbf{G}^{\prime\prime}\rightarrow\mathbf{G}^{\prime\prime}/V=\mathbf{G}^{\prime}_{0}$ is the natural isomorphism (the projection to the second factor). Thus, we get a Levi decomposition $\mathbf{G}^{\prime\prime}=V\rtimes\mathbf{G}^{\prime}_{0}$, which is compatible with our fixed Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$.
Next, fix a Levi decomposition $\mathbf{G}^{\prime}=V^{\prime}\rtimes^{\prime}\mathbf{G}^{\prime}_{0}$. It is uniquely determined by a section of $\mathbf{G}^{\prime}\rightarrow\mathbf{G}^{\prime}_{0}$, i.e. an injective morphism $s^{\prime}\colon\mathbf{G}^{\prime}_{0}\rightarrow\mathbf{G}^{\prime}$ such that $s^{\prime}\circ p_{\mathbf{G}^{\prime}}=1_{\mathbf{G}^{\prime}_{0}}$. Now abuse of notation denote by $s^{\prime}$ the composite $\mathbf{G}^{\prime}_{0}\xrightarrow{s^{\prime}}\mathbf{G}^{\prime}<\mathbf{G}^{\prime\prime}$. Then $s^{\prime}$ defines a Levi decomposition $\mathbf{G}^{\prime\prime}=V\rtimes^{\prime}\mathbf{G}_{0}^{\prime}$.
By the general theory of algebraic groups, any two Levi decompositions of $\mathbf{G}^{\prime\prime}$ differ from the conjugation by some $v_{0}\in V({\mathbb{Q}})$. Now, the conclusion follows from Proposition B.3.(i).∎
@@ B.3. Quotient by a normal subgroup and weak Mumford–Tate domains
with the following properties: (i) ${\mathscr{D}}/N$ is a Mumford–Tate domain in some classifying space of mixed Hodge structures (which must be of weight $-1$ and 00) and $\mathrm{MT}({\mathscr{D}}/N)=\mathbf{G}/N$; (ii) each fiber of $p_{N}$ is an $N({\mathbb{R}})^{+}$-orbit. It is clearly true that $p_{N}$ is semi-algebraic. More precisely, let $M_{-1}/N$ be the maximal quotient of $M_{-1}$ on which $N$ acts trivially, and let $M/N=M_{-1}/N\oplus{\mathbb{Q}}$. For each $x\in{\mathscr{D}}$, the composition
@@
defines some Hodge structure on $M/N$ with weight $-1$ on $M_{-1}/N$ and weight 00 on the quotient ${\mathbb{Q}}$. The Hodge numbers on $M/N$ defined by $h_{x}$ do not depend on the choice of $x$. Let ${\mathscr{M}}^{\prime}$ denote the corresponding classification space. Then $\bar{h}_{x}$ defined a point $\bar{x}$ in ${\mathscr{M}}^{\prime}$. The images of such $\bar{x}$ form a Mumford–Tate domain denoted by ${\mathscr{D}}/N$. Thus, we have the desired quotient map $p_{N}:{\mathscr{D}}{\longrightarrow}{\mathscr{D}}/N$ which is both complex analytic and real semi-algebraic.
Assume $\Gamma$ is an arithmetic subgroup of $\mathbf{G}({\mathbb{Q}})$. Then the quotient $\Gamma\backslash{\mathscr{D}}$ is an orbifold. Denote by $\Gamma_{/N}$ the image of $\Gamma$ under the quotient $\mathbf{G}\rightarrow\mathbf{G}/N$. Then the quotient $p_{N}$ induces
@@
We will pay special attention to the fibers of $p_{N}$ (and $[p_{N}]$). More generally, we define:
@@
In this definition, if $x$ is taken to be a Hodge generic point, i.e. $\mathrm{MT}(x)=\mathbf{G}$, then the weakly Mumford–Tate domain thus obtained is a fiber of $p_{N}$.
@@ C.1. Definition and basic property of Intermediate Jacobians
Let us look at the following example from geometry. Let $f\colon X\rightarrow S$ be a smooth projective morphism of relative dimension $d$ over a complex quasi-projective variety such that each fiber is irreducible. For each $n\geq 1$, the relative intermediate Jacobian of the (polarizable) VHS $R^{2n-1}f_{*}{\mathbb{Z}}(n)_{X}$ is called the $n$-th intermediate Jacobian of $X\rightarrow S$ and is denoted by ${\mathscr{J}}^{n}(X/S)$. If $S$ is a point and $X=X$, then we simply write ${\mathscr{J}}^{n}(X)$.
@@
The dual of $H^{1,0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\subseteq H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})$ gives $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})^{\vee}\rightarrow\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})$. Thus the de Rham–Betti comparison $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\cong H^{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{C}})$ gives rise to a Hodge filtration ${\mathscr{F}}_{{\mathscr{J}}}^{\bullet}$ on ${\mathscr{V}}$ defined by ${\mathscr{F}}_{{\mathscr{J}}}^{1}=0$, ${\mathscr{F}}_{{\mathscr{J}}}^{-1}={\mathscr{V}}$ and
@@
We computed in the proof of Lemma C.2 that ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{V}}_{s}/({\mathscr{F}}^{n}_{s}+{\mathbb{V}}_{{\mathbb{Z}},s})$. Hence $\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})={\mathscr{V}}_{s}/{\mathscr{F}}^{n}_{s}$. So $({\mathscr{F}}_{{\mathscr{J}}}^{0})_{s}={\mathscr{F}}^{n}_{s}$ by defintion. We are done. ∎

### v2

@@ B.2. Mumford–Tate domains
The unipotent radical of $\mathbf{G}=\mathrm{MT}({\mathscr{D}})$ equals $\mathbf{G}_{-1}:=V\cap\mathbf{G}$ by reason of weight. Let $\mathbf{G}_{0}:=\mathbf{G}/\mathbf{G}_{-1}$ be the reductive part. Set ${\mathscr{D}}_{0}:=p({\mathscr{D}})\subseteq{\mathscr{M}}_{0}$ for the map $p$ defined in (B.6). Then ${\mathscr{D}}_{0}$ is a $\mathbf{G}_{0}({\mathbb{R}})^{+}$-orbit and is in fact a (pure) Mumford–Tate domain in the classifying space ${\mathscr{M}}_{0}$, and $\mathrm{MT}_{x_{0}}<\mathbf{G}_{0}$ for all $x_{0}\in{\mathscr{D}}_{0}$.
@@
(B.8) $p\colon\mathbf{G}\rightarrow\mathbf{G}_{0}=\mathbf{G}/\mathbf{G}_{-1}\quad\text{ and }\quad p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}.$
Fix a Levi decomposition $\mathbf{G}=\mathbf{G}_{-1}\rtimes\mathbf{G}_{0}$, such as one given by B.7. Identify $\mathbf{G}_{0}$ with $\{0\}\times\mathbf{G}_{0}$.
Recall that each $x_{0}\in{\mathscr{M}}_{0}$ endows $V$ with a Hodge structure of weight $-1$. For $x_{0}\in{\mathscr{D}}_{0}$, $\mathbf{G}_{-1}$ is a sub-Hodge structure because $\mathbf{G}_{-1}$ is a $\mathbf{G}_{0}$-submodule of $V$ and that $\mathrm{MT}_{x_{0}}<\mathbf{G}_{0}$.
Finally, consider the constant bundle $\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0}\rightarrow{\mathscr{D}}_{0}$. Define the holomorphic subbundle $F^{0}(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})$ to be such that the fiber over each $x_{0}\in{\mathscr{D}}_{0}$ is $F^{0}_{x_{0}}\mathbf{G}_{-1,{\mathbb{C}}}$.
@@
    $\mathbf{G}_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}\xrightarrow{\sim}{\mathscr{D}},\quad(v,x_{0})\mapsto\mathrm{Int}(v)\circ h_{x_{0}}.$
@@
    The complex structure on ${\mathscr{D}}$ is given by ${\mathscr{D}}{\overset{\sim}{{\longrightarrow}}}(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})$.
@@
    (B.9) $\mathbf{G}_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0}\longrightarrow(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0}).$
@@
The Levi decomposition $\mathbf{G}=\mathbf{G}_{-1}\rtimes\mathbf{G}_{0}$ induces a bijection $\mathbf{G}({\mathbb{R}})^{+}x\simeq\mathbf{G}_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}$, which is semi-algebraic. This establishes (i).
Set ${\mathscr{X}}:=\mathbf{G}({\mathbb{R}})^{+}\mathbf{G}_{-1}({\mathbb{C}})\cdot h_{x}\subseteq{\mathrm{Hom}}({\mathbb{S}}_{{\mathbb{C}}},\mathbf{G}_{{\mathbb{C}}})$, where the action is given by conjugation. The quotient $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}$ induces a natural surjective map ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and by [Pin89, 1.8(a)] each fiber of this map is a $\mathbf{G}_{-1}({\mathbb{C}})$-torsor.
The Levi decomposition $\mathbf{G}=\mathbf{G}_{-1}\rtimes\mathbf{G}_{0}$ induces a global section of ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and hence an isomorphism ${\mathscr{X}}\simeq\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0}$ over ${\mathscr{D}}_{0}$.
@@
By [Pin89, 1.8(b)], for each $x\in{\mathscr{D}}$, the fiber $\varphi^{-1}(x)$ is a principle homogeneous space under $F^{0}_{x_{0}}\mathbf{G}_{-1,{\mathbb{C}}}$ with $x_{0}=p(x)$. Hence $\varphi$ gives a bijection $(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(\mathbf{G}_{-1}({\mathbb{C}})\times{\mathscr{D}}_{0})\rightarrow{\mathscr{D}}$ over ${\mathscr{D}}_{0}$. Moreover, the complex structure of ${\mathscr{D}}$ is precisely given by this bijection; see [GK24, Proof of Prop. 2.6 in Appendix A] which is a consequence of [Pea00, Thm. 3.13]. This establishes (ii).
@@
Let $\widetilde{Z}_{0}\subseteq{\mathscr{D}}_{0}$ be an irreducible algebraic subset. Then for any $a\in\mathbf{G}_{-1}({\mathbb{R}})$, the subset $\{a\}\times\widetilde{Z}_{0}\subseteq\mathbf{G}_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}\simeq{\mathscr{D}}$ is irreducible algebraic.
@@
$\{a\}\times\widetilde{Z}_{0}$ is clearly semi-algebraic in ${\mathscr{D}}$. In view of (B.9), $\{a\}\times\widetilde{Z}_{0}$ is also complex analytic in ${\mathscr{D}}$. Hence, we are done. ∎
@@
Let ${\mathscr{D}}^{\prime}$ be a sub-Mumford–Tate domain of ${\mathscr{D}}$. Then under the identification ${\mathscr{D}}=\mathbf{G}_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), we have ${\mathscr{D}}^{\prime}=(\mathbf{G}_{-1}^{\prime}({\mathbb{R}})+v_{0})\times p({\mathscr{D}}^{\prime})$ for some $v_{0}\in\mathbf{G}_{-1}({\mathbb{Q}})$.
@@
It suffices to apply (B.7) to both ${\mathrm{MT}}({\mathscr{D}})$ and ${\mathrm{MT}}({\mathscr{D}}^{\prime})$. ∎
@@ B.3. Quotient by a normal subgroup and weak Mumford–Tate domains
with the following properties: (i) ${\mathscr{D}}/N$ is a Mumford–Tate domain in some classifying space of mixed Hodge structures (which must be of weight $-1$ and 00) and $\mathrm{MT}({\mathscr{D}}/N)=\mathbf{G}/N$; (ii) each fiber of $p_{N}$ is an $N({\mathbb{R}})^{+}$-orbit. It is clearly true that $p_{N}$ is semi-algebraic. More precisely, let $V/N$ be the maximal quotient of $V$ on which $N$ acts trivially, and let $M/N=V/N\oplus{\mathbb{Q}}$. For each $x\in{\mathscr{D}}$, the composition
@@
defines some Hodge structure on $M/N$ with weight $-1$ on $V/N$ and weight 00 on the quotient ${\mathbb{Q}}$. The Hodge numbers on $M/N$ defined by $h_{x}$ do not depend on the choice of $x$. Let ${\mathscr{M}}^{\prime}$ denote the corresponding classification space. Then $\bar{h}_{x}$ defined a point $\bar{x}$ in ${\mathscr{M}}^{\prime}$. The images of such $\bar{x}$ form a Mumford–Tate domain denoted by ${\mathscr{D}}/N$. Thus, we have the desired quotient map $p_{N}:{\mathscr{D}}{\longrightarrow}{\mathscr{D}}/N$ which is both complex analytic and real semi-algebraic.
Assume $\Gamma$ is an arithmetic subgroup of $\mathbf{G}({\mathbb{Q}})$. Then, the quotient $\Gamma\backslash{\mathscr{D}}$ is an orbifold. Denote by $\Gamma_{/N}$ the image of $\Gamma$ under the quotient $\mathbf{G}\rightarrow\mathbf{G}/N$. Then the quotient $p_{N}$ induces
@@
We will pay special attention to the fibers of $p_{N}$ (and $[p_{N}]$). More generally, we define
@@
In this definition, if $x$ is taken to be a Hodge generic point, i.e. $\mathrm{MT}(x)=\mathbf{G}$, then the weak Mumford–Tate domain thus obtained is a fiber of $p_{N}$.
@@ C.1. Definition and basic property of Intermediate Jacobians
Let us look at the following example from geometry. Let $f\colon X\rightarrow S$ be a smooth projective morphism of relative dimension $d$ over a complex quasi-projective variety such that each fiber is irreducible. For each $n\geq 1$, the relative intermediate Jacobian of the (polarizable) VHS $R^{2n-1}f_{*}{\mathbb{Z}}(n)_{X}$ is called the $n$-th intermediate Jacobian of $X\rightarrow S$ and is denoted by ${\mathscr{J}}^{n}(X/S)$. If $S$ is a point, then we simply write ${\mathscr{J}}^{n}(X)$.
@@
The dual of $H^{1,0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\subseteq H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})$ gives $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})^{\vee}\rightarrow\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})$. Thus the de Rham–Betti comparison $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\simeq H^{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{C}})$ gives rise to a Hodge filtration ${\mathscr{F}}_{{\mathscr{J}}}^{\bullet}$ on ${\mathscr{V}}$ defined by ${\mathscr{F}}_{{\mathscr{J}}}^{1}=0$, ${\mathscr{F}}_{{\mathscr{J}}}^{-1}={\mathscr{V}}$ and
@@
We computed in the proof of Lemma C.2 that ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathbb{V}}_{{\mathbb{Z}},s}\backslash{\mathscr{V}}_{s}/{\mathscr{F}}^{n}_{s}$. Hence $\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})={\mathscr{V}}_{s}/{\mathscr{F}}^{n}_{s}$. So $({\mathscr{F}}_{{\mathscr{J}}}^{0})_{s}={\mathscr{F}}^{n}_{s}$ by defintion. We are done. ∎